# Promptlinecolor=black,linewidth=2pt, leftmargin=1cm,rightmargin=1cm MALLM-GAN: Multi-Agent Large Language Model as Generative Adversarial Network for Synthesizing Tabular Data

### v3

@@ 4.2 Performance Evaluation
We evaluate the performance of synthetic data generation models from two perspectives: Privacy leakage by Distance to Closest Records (DCR) and Machine Learning Efficiency (MLE) (10; 42).
MLE. To assess the utility of our synthetic data, we use it to train supervised model and test prediction accuracy on real data (). The Adult data is used for a classification task, while the other three datasets are used for regression. For classification, we fit logistic regression model, random forest, and Support Vector Machine model, XGBoost Classifier, calculating F1 score. For regression, we fit linear regression, random forest, XGBoost Regressor and calculate . For each model, we report the average of the best scores for each random seed. We also fit models using real data as a gold standard of the MLE for comparison. As a result, MALLM-GAN generated high-quality synthetic tabular data across multiple datasets and training data size, outperforming baselines (Table 2), specially with small training sizes (). This indicates MALLM-GAN’s robustness to smaller sample sizes, unlike baselines that require more data. MALLM-GAN also outperformed baselines on both public and private datasets, suggesting it does not rely on the pre-trained LLM’s memorization.
| Public dataset | Private dataset | |||||
| Adult () | Asia () | Insurance() | ATACH() | ERICH() | ||
| N=100 | Real data | 0.86 | 0.83 | 0.82 | 0.26 | -0.04 |
| SMOTE* | ||||||
| TabDDPM | - | |||||
| CTGAN | ||||||
| TVAE | ||||||
| Be-GReaT | ||||||
| MALLM-GAN | ||||||
| N=200 | Real data | 0.85 | 0.83 | 0.83 | 0.27 | 0.16 |
| SMOTE* | ||||||
| TabDDPM | - | |||||
| CTGAN | ||||||
| TVAE | ||||||
| BeGReaT | ||||||
| MALLM-GAN | ||||||
| N=400 | Real data | 0.83 | 0.84 | 0.85 | 0.31 | 0.18 |
| SMOTE* | ||||||
| TabDDPM | - | |||||
| CTGAN | ||||||
| TVAE | ||||||
| Be-GReaT | ||||||
| MALLM-GAN | ||||||
| N=800 | Real data | 0.71 | 0.84 | 0.85 | 0.40 | 0.21 |
| SMOTE* | ||||||
| TabDDPM | - | |||||
| CTGAN | ||||||
| TVAE | ||||||
| Be-GReaT | ||||||
| MALLM-GAN | ||||||
DCR distributions. The DCR metric assesses the realism and diversity of synthetic data. It determines whether synthetic data points are too similar to the real data points (potential privacy leakage) or too dissimilar (hurting the utility of the synthetic data). The DCR is defined as Low DCR indicates that synthetic data are very close to real data points, implying a privacy leakage, as synthetic data too closely mimic the real data. We chose to use -norm distance to measure the distance between two data points (6). For the categorical variables, the distance is 1 if two categories are different; otherwise, 0. As a result, MALLM-GAN achieved similar or higher DCR levels compared to baseline models (Fig. 2), implying effective privacy protection without compromising MLE. Overall, MALLM-GAN demonstrated superior performance in generating synthetic data with small data as balancing privacy and utility.
@@ 4.3 Ablation study
Number of example in in-context few shot learning. Due to the LLM’s limited context length, we implemented a“batches in a batch” method to leverage all training data within these constraints (Section 3.2.1). We varied the number of examples and found to be optimal, achieving high DCR without compromising MLE (Supplement Section 2.5).
Causal structure and Optimization. To assess the impact of each component on overall performance, we examined the contribution of the causal structure in the data generation process and the LLM as an optimizer. We compared the full model, which includes both components, to a version without them, similar to CLLM (31) without post-processing data selection (Table 3). The ablation study showed that incorporating the causal structure alone did not significantly improve the MLE compared to a model with only in-context few-shot learning. However, the LLM optimizer improved using prior knowledge encoded in LLM and finally achieved the highest MLE. Incorporating external knowledge into LLMs has been shown to significantly improve the quality of generated text, similar to retrieval-augmented generation (RAG) (20). Our approach shares this concept by incorporating a “knowledge” graph but optimizes the knowledge itself through adversarial optimization.
@@ 4.4 Optimization trajectory of data generation process
A key advantage of MALLM-GAN is its transparent data generation process, described in natural text, which allows us to observe the evolution trajectory of the data generation mechanism during adversarial optimization. We present examples of optimization trajectories. We showed how the causal structure evolves to ground truth (Fig. 3) over iteration. We used the Asia dataset because it has known ground-truth causal structures and reported graph edit distance (GED) between ground truth and identified causal structures. In this example, the heuristically initialized causal structure gradually converges to ground truth, thanks to the knowledge obtained from the pre-trained LLM. Different convergence patterns were observed with different initialization strategies (Table 9), supporting the benefit of our heuristic initialization. We also investigated how the task instruction in the generator prompt gets sophisticated and how the discriminator’s accuracy changes over iterations. In Table 4, the task instructions evolved to include specific details, and the discriminator’s accuracy decreased, implying that synthetic data gets indiscriminative to real data.
| Iteration | Task instruction | Score |
@@ 4.6 Limitations
The proposed framework has several shortcomings. Firstly, due to the limited context length of the LLM, our model struggles with high-dimensional datasets with too many categorical variables, which make the context information lengthy and reduce the success rate of data generation. Another limitation of synthetic data generation introduced by LLM is that the LLM struggles with random number generation as pointed out in (15), which cast negative effects on our framework’s potential when dealing with datasets of many continuous variables. Additionally, mimicking the traditional GAN framework, it suffers from a theoretical convergence guarantee. While our model performs well with small sample sizes, showing better results than other baselines, the improvement diminishes with larger datasets. Moreover, the training and generation process is costly when dealing with large data volumes.
@@ 5 Conclusion
Supplementary
1 Example of prompts and output
@@
2 Experiment details
2.1 Benchmark datasets descriptions
@@
| Adult | 32,561 | 14 | The dataset include people’s social economic factors and demographics with the label that indicates whether their income is higher than 50k. | (4) |
| Medical Insurance | 2,772 | 7 | This is a dataset used to describe the paitents’ demographics with their health insurance bills. | (1) |
@@ 2.2 Hyperparameters
2.3 Comparison among different kinds of discriminators
| N = 100 | N = 200 | N = 400 | N = 800 | ||
| Adult (F1 score) | XGBoost | ||||
| Logistic regression | |||||
@@ 2.3 Comparison among different kinds of discriminators
2.4 Computing resource details
The model proposed in this study does not require extensive computing resource for fine-tuning. However, this model require access to Azure service. For other baseline models, they are implemented on one NVIDIA H100 80GB HBM3 GPU.
2.5 Number of examples in in-context few shot learning
Given the limited context length that the LLM can understand, we proposed “batches in a batch” method to leverage all training data in limited context length in the generator LLM (Section 3.2.1). We varied the number of few-shot examples by and measured the MLE (Fig. 5) and DCR distribution (Table 7, 8) to find the optimal number . As a result, the increasing number of examples did not always increase the MLE of synthetic data (Fig. 5) but decreased the DCR (Table 7), thus increasing privacy concerns. Instead, achieved sufficiently high DCR without compromising MLE. The MLE did not increase with more examples because the more examples will increase the context length and the generator LLM overlook some key context information. On the other hand, the DCR decreased with more examples because the generator LLM is more likely to stick to copy the provided examples. Interesting, the increasing number of examples does not affect the DCR of synthetic data generated from public dataset (Adult, Insurance).
@@ 2.5 Number of examples in in-context few shot learning
2.6 Optimization trajectory: example on insurance dataset
As seen in Table 9, the causal structure does not converge to the ground truth after 5 epochs. When initialized with the ground truth, the causal structure maintains slight fluctuations. Another example of causal structure discovery on insurance dataset were presented in Supplementary 2.6.
| Epoch | No Initialization | Heuristic Initialization | Ground Truth Initialization |
| 0 | 16 | 5 | 0 |
| 1 | 4 | 1 | 0 |
| 2 | 4 | 1 | 0 |
| 3 | 3 | 0 | 0 |
| 4 | 4 | 0 | 2 |
Here is the example for the Insurance dataset. Initially, the causal structure derived from heuristics had no edges (Fig. 6). Over iterative optimization, a stable causal structure emerged. The task instructions evolved to include specific details (Table 10). Our objective was to reduce the discriminator’s accuracy, and the score decreased over iterations, thanks to the optimized causal structure and task instruction.
| Iteration | Task instruction | Score |
| Epoch 1 | “The task is to generate synthetic data that accurately mirrors these causal relationships. The data should include variables such as ’age’, ’sex’, ’bmi’, ’children’, ’smoker’, ’region’, and ’charges’. Each variable should influence the others as per the causal structure, creating a realistic representation of possible real-world scenarios.” | 80.95% |
| Epoch 2 | “Your task is to generate synthetic data that faithfully represents these causal relationships. The data should encompass variables such as ’age’, ’bmi’, ’children’, ’smoker’, ’region’, and ’charges’. Each variable should affect the others in accordance with the causal structure, providing a credible simulation of potential real-world scenarios” | 76.19% |
| Epoch 4 | “The ultimate goal is to generate synthetic data that accurately reflects these causal relationships. The synthetic data should incorporate factors such as ’age’, ’sex’, ’bmi’, ’children’, ’smoker’, ’region’, and ’charges’, and their influence on each other as indicated in the causal structure. The synthetic data should be convincing and representative of the real-world scenarios where these variables interact.” | 35.67% |

### v4

@@ 5.1 Performance Evaluation
We evaluate the performance of synthetic data generation models from two perspectives: Privacy leakage by Distance to Closest Records (DCR) and Machine Learning Efficiency (MLE) (10; 41).
MLE. To assess the utility of our synthetic data, we use it to train a supervised model and test prediction accuracy on real data (). The Adult data are used for a classification task, while the other three datasets are used for regression. For classification, we fit the logistic regression model, random forest, and Support Vector Machine model, XGBoost Classifier, calculating the score. For regression, we fit linear regression, random forest, XGBoost Regressor and calculate . For each model, we report the average of the best scores for each random seed. We also fit models using real data as a gold standard of the MLE for comparison. As a result, MALLM-GAN generated high-quality synthetic tabular data across multiple datasets and training data sizes, outperforming baselines (Table 2), especially with a high dimension setting (, e.g. ATACH2 and ERICH). This indicates MALLM-GAN‘s robustness to smaller sample sizes, unlike baselines that require more data. While TabPFN also achieves comparable performance and scales well with increasing sample sizes, it has notable limitations—its effectiveness declines when the number of categorical levels exceeds 10 (e.g. Adult and ERICH) or when all variables are categorical (e.g. Aisa), both of which are common scenarios in real-world datasets. Furthermore, MALLM-GAN outperformed the baselines in both public and private datasets, suggesting that it does not rely on the pre-trained LLM’s memorization.
We also benchmark our model on medium dataset scenario (N=400, 800). The results, as shown in Table 6, demonstrate that the data driven model can scale their performance as the dataset size increase, while our model can still achieve a comparable performance.
DCR distributions. The DCR metric assesses the realism and diversity of synthetic data. It determines whether the synthetic data points are too similar to the real data points (potential privacy leakage) or too dissimilar (hurting the utility of the synthetic data). The DCR is defined as Low DCR indicates that the synthetic data are very close to real data points, implying a privacy leakage, as synthetic data too closely mimic the real data. We used -norm distance to measure the distance between two data points (6). For the categorical variables, the distance is 1 if two categories are different; otherwise, 0. As a result, MALLM-GAN achieved similar or higher DCR levels compared to baseline models (Fig. 2, an example on data Adult with N=100; evaluation on other datasets are provided in Figure 6), implying effective privacy protection without compromising MLE. Overall, MALLM-GAN demonstrated superior performance in generating synthetic data with small data as balancing privacy and utility.
@@ 5.2 Ablation study
Number of example in in-context few shot learning. Due to the LLM’s limited context length, we implemented a "batches in a batch" method to leverage all training data within these constraints (Section 3.2.1). We varied the number of examples and found the optimal varied by different datasets considering their characteristics (Fig. 7). As a result, the increasing number of examples did not always increase the quality of synthetic data in terms of MLE and DCR (Fig. 7 and Table 10). The MLE did not increase with more examples because more examples will increase the context length, and the generator overlooks some key context information.
Causal structure and Optimization.
To assess the impact of each component on overall performance, we examined the contribution of the causal structure in the data generation process and the LLM as an optimizer. We compared the full model, which includes both components, to a version without them, similar to CLLM (33) without post-processing data selection (Table 3). The ablation study showed that incorporating the causal structure alone did not significantly improve the MLE compared to a model with only in-context few-shot learning. However, the LLM optimizer improved using prior knowledge encoded in LLM and finally achieved the highest MLE. Incorporating external knowledge into LLMs has significantly improved the quality of generated text, similar to retrieval-augmented generation (RAG) (23). Our approach shares this concept by incorporating a knowledge graph but optimizing the knowledge itself through adversarial optimization.
@@ 5.3 Optimization trajectory of data generation process
A key advantage of MALLM-GAN is its transparent data generation process, described in natural text, which allows us to observe the evolution trajectory of the data generation mechanism during adversarial optimization. We present examples of optimization trajectories. We showed how the causal structure evolves to the ground truth (Fig. 4) over iteration. We used the Asia dataset because it has known ground-truth causal structures and reported graph edit distance (GED) between ground truth and identified causal structures. In this example, the heuristically initialized causal structure and the causal structure even without initialization gradually converge to ground truth, thanks to the knowledge obtained from the pre-trained LLM. Different convergence patterns were observed with different initialization strategies (Fig. 3).
We also investigated how the task instruction in the generator prompt gets sophisticated and how the discriminator’s accuracy changes over iterations. In Table 4, the task instructions evolved to include specific details, and the discriminator’s accuracy decreased, implying that synthetic data become indiscriminate to real data.
| Iteration | Task instruction | Score ↓ |
@@ 5.5 Limitations
The proposed framework has several shortcomings. Firstly, due to the limited context length of the LLM, our model struggles with high-dimensional datasets with too many categorical variables, which makes the context information lengthy and reduces the success rate of data generation. Another limitation of synthetic data generation introduced by LLM is that the LLM struggles with random number generation as pointed out in (17), which casts negative effects on our framework’s potential when dealing with datasets of many continuous variables. Additionally, mimicking the traditional GAN framework, it suffers from a theoretical convergence guarantee. While our model performs well with small sample sizes, showing better results than other baselines, the improvement diminishes with larger datasets. Moreover, the training and generation process is costly when dealing with large data volumes. A demonstration of computational cost is shown in Table 12
@@ 6 Conclusion
Appendix A Example of prompts and output
@@
Appendix B Experiment details
B.1 Benchmark datasets descriptions
@@
| Adult | 32,561 | 14 | The dataset includes people’s socioeconomic factors and demographics, with the label that indicates whether their income is higher than 50k. | (3) |
| Magic | 19,020 | 10 | It is a simulated registration of high-energy gamma particles in a ground-based atmospheric Cherenkov gamma telescope using the imaging technique | (5) |
| Medical Insurance | 2,772 | 7 | This dataset describes the paitents’ demographics with their health insurance bills. | (1) |
@@ B.2 Hyperparameters
B.3 Additional Experiments on Medium Size Datasets
We also benchmark our model on the
| Public dataset | Private dataset | ||||||
| Adult () | Magic() | Asia () | Insurance() | ATACH() | ERICH() | ||
| N=400 | Real data | 0.83 | 0.82 | 0.84 | 0.85 | 0.31 | 0.18 |
| SMOTE* | |||||||
| TabDDPM | - | ||||||
| CTGAN | |||||||
| TVAE | |||||||
| Be-GReaT | |||||||
| Tabsyn | - | - | |||||
| TabPFN | - | - | - | ||||
| MALLM-GAN | |||||||
| N=800 | Real data | 0.71 | 0.81 | 0.84 | 0.85 | 0.40 | 0.21 |
| SMOTE* | |||||||
| TabDDPM | - | ||||||
| CTGAN | |||||||
| TVAE | |||||||
| Be-GReaT | |||||||
| Tabsyn | - | - | |||||
| TabPFN | - | - | - | - | |||
| MALLM-GAN | |||||||
B.4 DCR evaluation on other datasets
The following figures are to evaluate DCR on other datasets:
B.5 Comparison of different discriminators
In the study, we compare 3 different kinds of supervised classification models as the role of a discriminator. An experiment was conducted on the Adult dataset’s sub-sample to demonstrate the discriminator’s effects on the quality of the generated data.
| N = 100 | N = 200 | N = 400 | N = 800 | |
| XGBoost | ||||
| Logistic regression | ||||
@@ B.5 Comparison of different discriminators
B.6 Computing resource details
The model proposed in this study does not require extensive computing resources for fine-tuning. However, this model requires access to Azure service. For other baseline models, they are implemented on one NVIDIA A100 40GB GPU.
B.7 Ablations studies on number n of provided real samples
@@
| 1 | 2 | 3 | 4 | 5 | |
| Adult | 5, 6, 10 | 5, 7, 12 | 4, 6, 9 | 4, 6, 10 | 4, 6, 11 |
| Magic | 23, 29, 37 | 24, 33, 52 | 30, 44, 65 | 37, 53, 72 | 40, 55, 73 |
| ATACH2 | 61, 73, 88 | 72, 87, 97 | 69, 78, 89 | 66, 75, 94 | 67, 83, 104 |
| ERICH | 53, 61, 79 | 59, 78, 96 | 59, 74, 101 | 56, 72, 96 | 50, 64, 88 |
| 1 | 2 | 3 | 4 | 5 | |
| Adult | 4, 7, 10 | 5, 7, 11 | 5, 6, 10 | 4, 7, 10 | 4, 7, 11 |
| Magic | 45, 61, 87 | 44, 57, 82 | 45, 58, 88 | 46, 59, 86 | 45, 60, 86 |
| ATACH2 | 84, 100, 120 | 82, 99, 122 | 81, 97, 125 | 79, 98, 124 | 82, 103, 128 |
| ERICH | 70, 87, 110 | 66, 82, 111 | 51, 82, 104 | 62, 80, 108 | 62, 80, 117 |
Appendix C Cost Analysis
| Dataset | Sample Size | Number of Epochs | Number of Examples Per Call | Training Time | Inference Time |
| Asia | 50 | 10 | 4 | 13.5 | 5.9 |
| Asia | 200 | 4 | 4 | 24.8 | 3.8 |
| ERICH | 100 | 5 | 1 | 62.7 | 14.6 |
| ERICH | 200 | 4 | 1 | 92.5 | 23.7 |